OpenStudy (anonymous):

If f(x) = x2 + 2, find f(4). 18 10 6 2

OpenStudy (anonymous):

slyethrin?

OpenStudy (anonymous):

HA! Ravenclaw?

OpenStudy (anonymous):

\[ f(x) = x^2 + 2\\\\f(4)=4^2 +2=18 \]

OpenStudy (anonymous):

nope slyethrin..just liked luna's name

OpenStudy (anonymous):

^ he's right

OpenStudy (anonymous):

:D